### v1

# A solvable embedding mechanism for one-dimensional spinless and Majorana fermions in higher-dimensional spin-$\frac{1}{2}$ magnets

Abstract
We write down a class of two-dimensional quantum spin-$\frac{1}{2}$ Hamiltonians whose eigenspectra are exactly solvable via the Jordan-Wigner transformation. The general structure corresponds to a suitable grid composed of $XY$ or $XX$-Ising spin chains and $ZZ$-Ising spin chains and is generalizable to higher dimensions. They can host stacks of one-dimensional spinless fermion liquids with gapless excitations and power-law correlations coexisting with ordered spin moments (localized spinless fermions). Bond-dependent couplings thus can be an alternate mechanism than geometric frustration of $SU(2)$-symmetric couplings to obtain spinless fermionic excitations. Put in a different way, bond-dependent couplings allow for an embedding of one-dimensional spinless fermion (Tomonoga-Luttinger) liquids and solids and also Majorana excitations in higher dimensions. They can accommodate a simpler set of site-local conserved quantities apart from the more intricate, interlocked set of plaquette-local or bond-local conserved quantities in Kitaev’s honeycomb model with Majorana excitations. The proposed grid structure may provide an architecture for quantum engineering with controllable qubits.
Introduction:– Emergent fractionalized excitations are a remarkable manifestation of strongly correlated many-body physics. A canonical example is the one-dimensional ($1d$) spin-$\frac{1}{2}$ antiferromagnetic chain with fermionic spinon excitations which are colloquially said to be “half” of regular (bosonic) spin-wave excitations or magnons in presence of magnetic ordering. Magnons being close to product quantum states with low quantum entanglement allow one to form a classical image as quanta of smooth undulations of the ordered magnetic moments like phonons for lattice vibrations and photons for electromagnetic waves. Spinons on the other hand escape simple classical images. They are rather a complicated superposition of the underlying spin states with extremely high quantum entanglement. Such emergent excitations are thus fascinating apart from the qualitative new effects they engender [1].
Finding solvable models where such physics is explicit is of importance as proofs-of-principle. They serve as counterparts to field theoretic arguments that rely on low-energy approximations, perturbative renormalization group treatments and universality [2]. The downside is that solvability is often restricted to points in the phase diagram [3]. The Kitaev honeycomb model is a famous solvable model of fractionalized (free fermionic) Majorana excitations [4, 5, 6] on top of a spin liquid (SL) ground state in $2d$ $S=\frac{1}{2}$ antiferromagnets. In this work, we will write down a set of exactly solvable spin-$\frac{1}{2}$ models with bond-dependent couplings which host spinless fermionic or Majorana excitations akin to the Kitaev model but with rather different physics.
[…]
Figure 1: The grey dots, black dashed lines and the light red line are the $S=\frac{1}{2}$ degrees, $XY$ couplings and the Jordan-Wigner string respectively.
The green-highlighted bond in Fig. 1 has oppositely signed hopping amplitude compared to the rest of the bonds. However this change of sign in hopping amplitudes is a pure gauge artifact in $1d$. One can remove it by the following transformation: $c_{i}\rightarrow d_{i}$ on sites left of the green bond, and $c_{i}\rightarrow-d_{i}$ to the right.
Figure 2: Left panel shows the lowest energies (per site) from each conserved off-chain spin sector for an $L=8+8$ chain [8] with $\frac{J_{ZZ}}{J_{XY}}=0.5$. The $x$-axis indexes the different sectors through the decimal equivalent of the binary-valued off-chain spin configuration. One can verify that the two lowest energies corresponds to the two antiferromagnetic configurations for the off-chain spins. Middle panel shows the finite gap between the lowest energies from other sectors and the ground state energies for different system sizes. Right panel shows how this sector gap evolves with $\frac{J_{ZZ}}{J_{XY}}$ for $L=40$ chain.
This is special to $1d$ and hopping amplitude signs can not be “gauged away” in $d>1$ [9]. Thus for any $|\sigma\rangle=a|\uparrow\rangle+b|\downarrow\rangle$ for the disconnected spin, the solution is $a|\uparrow\rangle\otimes|\text{JWSL}_{c}\rangle+b|\downarrow\rangle\otimes|\text{JWSL}_{d}\rangle$. (Both $|\text{JWSL}_{c}\rangle$ and $|\text{JWSL}_{d}\rangle$ are degenerate can be linearly combined freely.)
Of course, one may just not include the disconnected spin in the JW-string, i.e. put it at the “very end” [10] to not affect the JW-hopping amplitudes, which gives $|\sigma\rangle\otimes|JWSL_{c}\rangle$ [11]. Thus off-chain spins can bring in only pure gauge artifacts if they are themselves not involved kinetically with the spin chain. Once off-chain spins are involved kinetically, then JW-solvability generally breaks down. The class of models that we write down here avoid such “off-chain kinetics”. The model construction will consist of collection of spin chains with on-chain kinetic terms and off-chain spins that are restricted to have (“interaction”) $ZZ$-Ising couplings with the on-chain spins. For such a bond-dependent scenario, the JW-transformation keeps working in obtaining the eigenspectra. Gauge artifacts of the sort above are removed by a judicious “snake-like” choice for the JW-string (Fig 3 left). The $Z$-component of the off-chain spins also become conserved quantities in the process that facilitates the solving for the full eigenspectrum. This leads to a $2d$ grid of $XY$/$XX$-Ising and $ZZ$-Ising chains [12]. Bond-dependent couplings in spin-$\frac{1}{2}$ Hamiltonians have assumed importance since Jackeli and Khaliullin’s mechanism [13]. The proposed class of bond-dependent models can thus form a new material class apart from so-called Kitaev materials [14, 15, 16, 17, 18, 19]. The models naturally divide into geometrically unfrustrated lattices containing no triangular motifs, and geometrically frustrated lattices containing triangular motifs [20]. A further sub-division concerns if the on-chain terms are of $XY$ or $XX$ type. In the following, we will consider ferromagnetic signs for the couplings without any loss of generality. There are local unitaries that can give all possible sign combinations and resultant ground states.
Figure 3: Example of a lattice without triangular motifs. The ground state sector corresponds to an antiferromagnetic (left), paramagnetic (right) arrangement of the conserved off-chain spins with gapped spinon on-chain excitations [21] Black and blue dashed lines represent $XY$ (left)/$XX$ (right) and $ZZ$-Ising couplings respectively. The light red line is the snake-like choice for the Jordan-Wigner string.
Figure 4: Left panel shows the lowest energies (per site) from each conserved off-chain spin sector for an $L=8+8$ chain [8] with $\frac{J_{ZZ}}{J_{XY}}=0.5$. The $x$-axis indexes the different sectors through the decimal equivalent of the binary-valued off-chain spin configuration. One can verify that the two lowest energies corresponds to the two ferromagnetic configurations for the off-chain spins. Middle panel shows the finite gap between the lowest energies from other sectors and the ground state energies for different system sizes. Right panel shows how this sector gap evolves with $\frac{J_{ZZ}}{J_{XY}}$ for these system sizes.
Unfrustrated $XY$:– An example of this construction is shown in the left panel of Fig. 3. The ground state is an antiferromagnet in the $Z$-direction. The JW-spectrum on the chains is gapped implying exponentially decaying transverse $XY$ correlations as well. The ground state can also be thought of as a spinless fermion charge density wave. In contrast to the Kitaev model where one takes recourse to Lieb’s theorem [22] to rigorously pin down the ground state sector, here we are not aware of any rigorous way of proving the ground state sector selection. Numerically [8] it becomes amply evident that the ground state sector indeed corresponds to an antiferromagnetic arrangement of the conserved off-chain spins (sketched in Fig. 3 left)  [21]. Such an arrangement corresponds to chemical potentials with alternating sign, i.e. modulating at wavevector $\pi$, for the on-chain JW-fermions which leads to gap opening near the Fermi level (at ±$\pm\pi/2$) along with charge density wave ordering at the same wavevector. Ground state sector selection can be argued for perturbatively in the $\frac{J_{XY}}{J_{ZZ}}\gg 1$ and $\frac{J_{XY}}{J_{ZZ}}\ll 1$ limits. For $\frac{J_{XY}}{J_{ZZ}}\ll 1$, the on-chain spin aligns with its partner off-chain spin to satisfy the ferromagnetic $ZZ$-Ising coupling. The neighboring on-chain spins now prefer to anti-align to gain energy through virtual hops. A different argument can also be made for this sector selection in the other $\frac{J_{XY}}{J_{ZZ}}\gg 1$ limit  [23]. The ground state interpolates smoothly between the two limits with the ratio of gap opening size to spinless Fermi sea bandwidth being controlled by $\frac{J_{XY}}{J_{ZZ}}$. Upon monitoring the lowest energy states in other conserved off-chain spin sectors, one finds a finite energy gap as well [24]. See Fig. 2. Thus this construction leads to gapped antiferromagnetic ground states [25]. A rigorous proof of this fact is desirable.
Frustrated $XY$ and $XX$:– An example of this construction is shown in Fig. 5 left whose ground state is a ferromagnet in the $Z$-direction [26]. The JW-fermion spectrum on the chains now remains gapless implying algebraically decaying transverse $XY$ correlations. Again the above is determined numerically [8]. Here too, we can make similar perturbative arguments in $\frac{J_{XY}}{J_{ZZ}}\gg 1$ and $\frac{J_{XY}}{J_{ZZ}}\ll 1$ limits for the ground state sector selection. There is an interpolation between the fully polarized state for $\frac{J_{ZZ}}{J_{XY}}>1$ and a half-filled spinless Fermi sea as $\frac{J_{ZZ}}{J_{XY}}\rightarrow 0$. Upon monitoring the lowest energy states in other conserved off-chain spin sectors, one again finds a finite energy gap as shown in Fig. 4 [27]. Thus this construction leads to ferromagnetic ground states with gapless on-chain excitations. A rigorous proof of the ground state sector selection is again desirable. The $XX$-Ising variant of this construction has related physics (Fig. 5 right). The ground state corresponds to a ferromagnetic configuration of the off-chain spins which effectively provides a transverse field to the on-chain spins. Thus depending on $\frac{J_{ZZ}}{J_{XX}}$, an Ising ordered or disordered state obtains for the on-chain spins which in terms of the JW fermions would be a topologically non-trivial or trivial superconducting state respectively with the transition at $\frac{J_{ZZ}}{J_{XX}}=0.5$ on this (kagome) lattice. The sector gaps are also finite similar to the $XY$ variant.
Figure 5: Example of a lattice with triangular motifs. The ground state sector corresponds to a ferromagnetic arrangement of the conserved off-chain spins with gapless spinon on-chain excitations. Black and blue dashed lines represent $XY$ (left)/$XX$ (right) and $ZZ$-Ising couplings respectively.
Unfrustrated $XX$:– This last case is the most unusual and interesting. It is the $XX$-Ising variant on lattices without triangular motifs. The ground state here turns out to be an Ising ordered/disordered ground state for the on-chain spins coexisting with a (classical) paramagnetic state for the off-chain spins! In other words, all the conserved off-chain spin sectors have the same spectrum as sketched in Fig. 3. This property arises due to an additional structure present in this case that was absent in the previous cases. This is due to the presence of additional conserved quantities apart from the off-chain spins that are mutually anticommuting [28]. The additional conserved quantities are products of $X$-component operators from sites on each of the Ising segments (blue dashed lines in Fig. 3). Since there are an extensive number of these operators coming from all the Ising segments, this leads to the extensive degeneracy associated with the off-chain spin configurations [29]. This seems pathological and in violation of the third law of thermodynamics [30]. Actually, this will rather be the physical state of affairs for generic temperature scales [31] similar to spin ices [32] and the SYK model [33]. In a “realistic” situation, other (smaller) couplings will then select the “true” ground state configuration for the off-chain spins to accord with the third law of thermodynamics. A natural guess in such a situation is to presume the existence of a (possibly gapless) “slow” mode primarily composed of the off-chain spins as excitations on top of the ground state, apart from the gapped Majorana excitations from the on-chain spins which will be higher in energy compared to the presumed slow mode [34]. The above structure also implies that the absence of triangular motifs is not an absolute requirement for this physics; rather one only requires $ZZ$-Ising (closed or open) segments with at least one conserved spin in the spirit of the above constructions connected through solely $XX$-Ising (or solely $YY$-Ising) segments.
Another physical consequence will be a two-step entropy release similar to that in the Kitaev spin liquids in presence of thermal fluctuations  [35, 36]. The first step would correspond to the entropy release due to the off-chain spins and the second step to the on-chain spins. The relative magnitude of the two steps will depend on the ratio of the number of off-chain spins to on-chain spins with it being 1:2 for Fig. 3. Thus these states can be considered as simpler variants of the Kitaev spin liquid. As differences, the conserved off-chain spins are geometrically independent and site-local in our constructions, whereas in the Kitaev model, the conserved plaquette operators are “interlocked” with common sites. The excitations also have a stripy nature in all of the above models, whereas the Majorana excitations in the Kitaev spin liquid are fully two-dimensional in nature. If there are additional conserved sites on the on-chain $XX$-Ising spin chains, then the resultant models are rendered effectively classical. The models considered here can therefore be said to be “in between” effectively classical states and fully quantum spin liquids as shown in Fig. 6.
Figure 6: A top-to-bottom sequence of models with increasing levels of quantum behavior. Black, red and green lines indicate $XX$,$YY$,$ZZ$-Ising couplings respectively. Left sequence has the model of Fig. 3 right at the middle. Right sequence has the Kitaev model at the bottom. The top effectively classical model in this sequence was studied by Mizoguchi [37]. Lieb lattice variants of the top row of brickwork lattices can also be envisaged.
[…]
It is also important to consider the stability of the grounds states discussed here with respect to natural deformations of the Hamiltonian, i.e. off-chain $XY$ couplings and on-chain $ZZ$ couplings. For the $XY$ variants with on-chain $ZZ$ couplings as deformations, one can invoke Luttinger liquid arguments to argue for stability [39, 2]. The unfrustrated $XY$ case would be stabler to on-chain $ZZ$ deformations due to the presence of the gap in the on-chain spinless fermionic excitations. The same would be expected for the frustrated and unfrustrated $XX$ case due to the gap in the on-chain spectrum in presence of Ising order or disorder.
Stability to off-chain deformations in the first three cases of unfrustrated $XY$, frustrated $XY$ and $XX$ models is evidenced by the presence of the sector gap as shown in Figs. 24. This is because it is precisely these deformations such as off-chain $XY$ or $XX$ couplings that would lead to quantum fluctuations in the off-chain spin $Z$-values that were otherwise conserved by construction. Presence of the sector gap implies that if the energy scale associated with these fluctuations is lesser than the gap, then the ground states should be stable till they become large enough. Thus the magnetic states discussed in this paper are stable states of magnetic matter. In the case of the unfrustrated $XX$ model with an extensive degeneracy from the paramagnetic off-chain spins, the off-chain deformations would be a possible mechanism for the development of the conjectured slow mode.
Figure 7: Some imagined quantum architectures.
Physical realization:– With a view towards physical realization, we can make the following remarks. In terms of materials, one may imagine an insulator material with localized $S=\frac{1}{2}$ moments with a crystal structure and quantum chemistry that accommodates a grid structure of easy-axis and easy-plane spin chains as the case may be. If the frustrated $XY$ case could be so realized, it would allow for highly anisotropic spin transport along the spin chains which can have practical applications such as in spin-based classical technologies. This could be an interesting avenue of exploration for the materials design and search community as alternative to the Kitaev material class with bond-dependent magnetic couplings.
But perhaps the more natural application of the ideas laid out in this work may be in the realm of quantum technologies involving artificial platforms with controllable qubits. The grid structure proposed here can provide a blueprint for a bus architecture for quantum information traffic. The conserved off-chain spins can sequester the spin information to the on-chain spin chains. Some imagined architectures are shown in Fig. 7. For example the right panel of Fig. 7 could be an use case where turning on and off the dashed $ZZ$-couplings would act as a valve for spin information flow. Interestingly, such an architecture has been recently reported in Ref. [40] with a different application as its focus.
Acknowledgements:– Funding support from SERB-DST, India via Grant No. MTR/2022/000386 and partially by Grant No. CRG/2021/003024 is acknowledged. I thank Sumilan Banerjee, Ganapathy Baskaran, Soumya Bera, Subhro Bhattacharjee, Hitesh Changlani, Kedar Damle, Nisheeta Desai, Giniyat Khaliullin, Gregoire Misguich, Laura Messio, Subroto Mukherjee, Ganpathy Murthy, Hridis Pal and Sumathi Rao for discussions. I thank Subhro in particular for the perturbative argument in the $J_{ZZ}/J_{XY}\ll 1$ limit. I am grateful to the Toulouse strongly correlated electron systems group, Laboratory of Theoretical Physics, University Paul Sabatier for the hospitality and discussions during a part of the work. I especially thank Fabien Alet for encouragement and feedback. I acknowledge the hospitality and support of ICTP, Trieste during a part of the work. I thank CEPIFRA for support during a part of the work.